40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 1 15:33:28 2022, Koji, Summary, BHD, Optomechanics configuration for the in-vacuum aux small optics Invac_opto_mechanics_v2.pdf
    Reply  Wed Feb 2 18:55:58 2022, Koji, Summary, BHD, Optomechanics configuration for the in-vacuum aux small optics Screen_Shot_2022-02-02_at_6.46.46_PM.png
       Reply  Wed Feb 2 21:16:23 2022, Koji, Summary, BHD, Optomechanics configuration for the in-vacuum aux small optics PXL_20220203_050156031.jpg
       Reply  Thu Apr 28 11:53:10 2022, Anchal, Summary, BHD, POP_SM4 and POP_SM5 Assembly 
          Reply  Tue May 3 14:02:09 2022, Anchal, Summary, BHD, Installed POP path in ITMX Chamber 
    Reply  Thu Apr 14 10:18:35 2022, Paco, Summary, BHD, 2 in oplev mirrors incompatible with LMR2V 
       Reply  Thu Apr 14 18:34:51 2022, Paco, Summary, BHD, ITMY Oplev reinstalled (Re: 2 in oplev mirrors incompatible with LMR2V) 
Message ID: 16826     Entry time: Tue May 3 14:02:09 2022     In reply to: 16817
Author: Anchal 
Type: Summary 
Category: BHD 
Subject: Installed POP path in ITMX Chamber 

[Anchal, JC]

I installed POP_SM4 and POP_SM5 in the ITMX chamber in the nominal positions. This must have affected the ITMX Oplev because I could see that one of the ITMX oplev beam was going through POP_SM5. It needs to be changed in order to follow the original plan. However, since POP_SM5 is a 1064 line mirror, it is transparent to the opleve beam, so maybe we can just use the ITMX oplev in the current fashion.

Next steps:

  • Get flashing back on the XARM.
  • Try to get the correct phase angle in POX11 so that we can lock XARM with IR too.
  • Inspect ITMX Oplev. The quadrant sum is low so maybe it needs adjustment in the in-ar table.
  • Check if ITMX oplev path needs to be adjusted inside the chamber.
ELOG V3.1.3-